OpenStudy (anonymous):

1. Which is a solution of the equation? 2x + y = 5 - I need someone to check and give me answers for this pre test! The test is similar but not the same. I like using answered pre tests as a study guide. Thanks!

OpenStudy (anonymous):

My answer is: (3,-1) Is this correct?

OpenStudy (anonymous):

looks good to me :)

OpenStudy (anonymous):

Okay thank you!! What is the slope of the line that passes through the points (3, 1) and (–2, 5)?

OpenStudy (anonymous):

wait the first one is not corect i think let me check

OpenStudy (anonymous):

it's 2 * 2 + 1 = 5

OpenStudy (texaschic101):

(3,-1) is a correct answer

OpenStudy (anonymous):

yea both are ok

OpenStudy (anonymous):

ok thank you

OpenStudy (texaschic101):

you need the slope ?

OpenStudy (anonymous):

yes

OpenStudy (texaschic101):

slope = (y2 - y1) / (x2 - x1) (3,1)..x1 = 3 and y1 = 1 (-2,5)...x2 = -2 and y2 = 5 now we sub slope = (5 - 1) / (-2 - 3) slope = 4/-5......so the slope is -4/5

OpenStudy (anonymous):

thank you. :) What about this one. What is the slope of a line parallel to the line whose equation is 3x – 4y = 8?

OpenStudy (anonymous):

and if you are seeing the question marks too, they are a minus sign lol

OpenStudy (texaschic101):

first we need to find the slope of the equation given because a parallel line will have the same slope. 3x - 4y = 8 -- put this in y = mx + b form, and m is your slope and b is your y intercept -4y = -3x + 8 y = 3/4x - 2 so our slope(m) = 3/4 a parallel line will have the same slope. A parallel line will not have the same y intercept. so your parallel equation, when given no points to work with, is y = 3/4x (+ or - any numbers besides -2) does that make sense to you ?

OpenStudy (texaschic101):

or you can do this.. slope will have to be 3/4 In y = 3/4x - 2......to find x intercept, sub in 0 for y 0 = 3/4x - 2 3/4x = -2 x = -2/(3/4) x = -2 * 4/3 x = -8/3....so your x intercept is (-8/3,0) y = mx + b slope(m) = 3/4 (-8/3,0)....x = -8/3 and y = 0 now sub 0 = 3/4(-8/3) + b 0 = -2 + b 2 = b so a parallel line would be : y = 3/4x + 2

OpenStudy (anonymous):

Okay thanks. So my answer is 3/4. :)

OpenStudy (anonymous):

What is the equation of the line in slope-intercept form? 4x - 9y =-36 how would i solve this?